• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ramesh

    Log entry time 07:58:10 on March19,2005

    Entry number 141905

    keyword=Shift Summary

    00:00 Shift started with ROC6 rebooting. I noticed the ROC5 failure comments popping up in the CODA during the run 2784. I tried to end the run, it failed. Rebooted the ROC6. In the meantime I also saw the ND/BB Scaler-window dying. I ended up with resetting SRC Trigger Supervisor, which brought the Scalers GUI back.


    Swing shift actually ended with the premature and unsuccessful run-end of 2784 instead of run 2783 (Simon's correction to his Shift summay).

    00:11-00:53 Run 2785(200k)
    00:53-01:39 Run 2786(200k)
    01:39-02:13 Run 2787(92k)

    02:13 MCC called saying they lost the laser-lock and they are approaching to gun-people; beam would be off for sometime (no time estimate).

    03:52 MCC gave a call saying that the "laser-lock" problem can not be solved at least in two/three hours. As an alternative, they will try to send us beam using Hall C laser and in another call during they said that the maximum possible beam will be only 20uA.

    Run 2788 (cosmic)

    04:08 20uA Beam back using Hall C laser.

    04:10-04:48 Run 2789(200k)
    04:48-05:25 Run 2790(200k)
    05:26-06:06 Run 2791(200k)
    06:07-06:44 Run 2792(200k)
    06:44-07:23 Run 2793(200k)
    07:23- Run 2794 passed to day shift.